OpenStudy (anonymous):

Please help A grease gun holds a cylindrical tube of grease that is 14 in. long and has a radius of 1.5 in. With each squeeze, 15 in3 of grease is used. How many times can the grease gun be squeezed before it is empty? Use 3.14 to approximate pi, and express your answer as a whole number.

OpenStudy (anonymous):

V=3.14r^2h <<< formula i know

OpenStudy (anonymous):

@sleepyhead314 @dan815 @Data_LG2 @iGreen.

OpenStudy (anonymous):

@sleepyhead314 @iGreen. @Data_LG2

OpenStudy (anonymous):

I'm not 100% sure if I'm right. So let x be the number of squeezes and y will be the volume I'll use y=mx+b where m will be the rate, which is 15 in^3 per squeeze and b will be the original volume so it will be \(\sf 3.14(1.5)^2 (14)\) Put these information together, you'll have \(\sf y= -15x+3.14(1.5)^2 (14)\) I put -15 because we have to subtract the volume that is being removed. Now, the question is asking for x when y=0 using the equation you can easily solve it.

OpenStudy (anonymous):

Find the volume

OpenStudy (anonymous):

\(V = \pi r^2 h\) Plug in what we know: \(V = (3.14)(1.5^2)(14)\) Multiply

OpenStudy (anonymous):

Actually, @Data_LG2 we can find the volume and divide by 15

OpenStudy (anonymous):

sorry guys computer went nuts

OpenStudy (anonymous):

oh right, idk know i keep making things so complicated.

OpenStudy (anonymous):

iGreen that equals 98.91

OpenStudy (anonymous):

**why ... os is lagging

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Yes, so divide by 15.

OpenStudy (anonymous):

sorry guys my computer isnt working @iGreen. did i get it right?

OpenStudy (anonymous):

Yeah, OS is lagging

OpenStudy (anonymous):

its 6.594

OpenStudy (anonymous):

Yes, so we can squeeze it 7 times before it's empty..

OpenStudy (anonymous):

so its 7

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

oh ok thx i have one more question ill open another tab and tag you